Let P = price, MR = marginal revenue, MC = marginal cost, and ATC = average total cost. In monopolistic competition, which of the following most accurately describes the long-run equilibrium conditions for a firm?

Group of answer choices

P > ATC, MR = MC, and P > MC

P > ATC, MR > MC, and P = MC

P = ATC, MR = MC, and P > MC

P = ATC, MR = MC, and P = MC

P = ATC, MR > MC, and P > MC
Answer #1

Answer

The correct answer is (c) P = ATC, MR = MC, and P > MC.

In order to maximize profit a firm produces that quantity at which MR = MC.

In the Long run when firms incur positive profit then new firms start entering which results in more number of substitutes available for the good and hence Residual demand faced by a firm decreases and this will continue till each firm making 0 profit i.e. Profit = 0.

Profit = TR - TC. Thus Profit = 0 => Tr - TC = 0 => TR = TC

=> P*Q = ATC*Q => P = ATC.

where TR = Total Revenue and TC = Total Cost and Q = quantity

Hence In the long run we have MR = MC and P = ATC.

Demand faced by a Monopolistic competition in downward sloping and because of this P is greater than MR and hence P > MR.

Thus we have P = ATC , MR = MC and P > MR => P > MR = MC.

Thus, In the Long run we have, P = ATC , MR = MC and P > MC.

Hence, the correct answer is (c) P = ATC, MR = MC, and P > MC.